OpenStudy (anonymous):

Help please : A total of $18,000 is invested at 3.25% and 2% simple interest. If yearly interest is $465, how much of the $18,000 is invested at each rate?

OpenStudy (kropot72):

Let a be the amount invested at 3.25% and let b be the amount invested at 2%. Now we can form two equations as follows: a + b = 18000 ....................(1) 0.0325a + 0.02b = 465 ......(2) You can find b by subtracting a from both sides of (1) giving: b = 18000 - a .....................(3) If you plug the value of b given in (3) into (2) you can solve to find a. Can you do that?

OpenStudy (anonymous):

yea im think im getting it

OpenStudy (anonymous):

so pretty much .0325a+0.02b=465 a+b=18000 and use elimination

OpenStudy (kropot72):

If you substitute the value for b given in (3) into (2) you get the following: 0.0325a + (0.02 * 18000) -0.02a = 465 ...........(4) Now can you solve equation (4) to find the value of a?

OpenStudy (anonymous):

uhm yea i got it a=8400

OpenStudy (anonymous):

thanks for your help :D

OpenStudy (anonymous):

greatly appreciated<3

OpenStudy (kropot72):

Your result for a is correct. Good work :)